# Adaptive Data Collection for Latin-American Community-sourced Evaluation of Stereotypes (LACES)

### v1

@@ 5.2 Unique Concepts
To quantify the conceptual novelty of the LACES dataset, we conducted an analysis to measure its thematic overlap with the CrowS-Pairs (Nangia et al., 2020), SeeGULL (Jha et al., 2023) and BBQ (Parrish et al., 2022) datasets.
First, we established a threshold for thematic similarity. This was accomplished by having four annotators manually label a set of 256 data points, marking pairs as either thematically related or distinct. Following this, we generated sentence embeddings for each data point in this annotated set. By correlating the cosine similarity scores of the embeddings with our manual labels, we calibrated and identified a threshold score for considering two concepts thematically equivalent.
Figure 5: Unique Concepts per Dataset. The bar chart displays the percentage of unique concepts for four distinct datasets: CrowS-Pairs, BBQ, SeeGULL, and LACES. The percentages are calculated based on a thematic similarity threshold.
Using this empirically determined threshold, a concept was considered present in another dataset if at least one data point in another dataset had an embedding with a similarity score above our threshold. As shown in Figure 5, the results demonstrate the percentage of unique concepts within each dataset when analyzed for thematic overlap, revealing that the LACES dataset contains the highest percentage of unique concepts at 55.60%. This is a notable increase compared to SeeGULL at 42.06% and a more significant increase compared to BBQ and CrowS-Pairs at 37.62% and 33.05% respectively.
Figure 6 provides a visualization of the conceptual overlap among the datasets. The width of the chords connecting the nodes quantifies the number of shared concepts between any two datasets, while the color of the chord indicates the dataset where the concept originated. The most notable insight is that the LACES dataset is the primary source of unique concepts, indicated by its large unconnected arc segment and the least overlap with other datasets. Furthermore, the BBQ and CrowS-Pairs datasets exhibit the largest mutual conceptual overlap, as demonstrated by the thickest connecting chord, suggesting they address the most similar set of underlying concepts. Interestingly, not all conceptual overlaps are symmetric; for example, a large number of concepts from BBQ can be found in SeeGULL but less than half that amount can be found the other way round.
@@ 5.3 Self-Debiasing
We examine whether a zero-shot self-debiasing technique proposed in recent work (Gallegos et al., 2025) in a non Latin American context generalizes to the socioculturally distinct content in LACES. We adopt the BBQ dataset and evaluation protocol (Parrish et al., 2022), instantiate it with LACES-derived questions, and compare the outcomes with the previously reported BBQ results. Each item asks in a multiple choice format which country is associated with a stereotypical attribute; an Unknown option is always available. The protocol details are provided in Appendix B. We evaluate three approaches for self-debiasing from Gallegos et al. (2025): baseline (B), explanation (E), and reprompting (R). The computed BBQ bias score is $\textsc{Bias}=(1-\textsc{ACC})\left[2\frac{n_{\textsc{biased}}}{m}-1\right]$, where ACC is accuracy over all items, $m$ is the number of non–Unknown predictions, and $n_{\textsc{biased}}$ counts the predictions that match the bias target. Table 2 shows mean and 95% confidence intervals (CI) obtained via 1,000 bootstrap replications.
GPT-3.5-Turbo GPT-4o-mini Llama 3.1-8B
BBQ dataset
B 0.12 [0.09, 0.15] 0.14 0.10
E 0.03 [0.01, 0.05] 0.01 0.00
R 0.04 [0.02, 0.06] 0.01 0.02
LACES dataset
B 0.53 [0.51, 0.55] 0.68 [0.66, 0.70] 0.55 [0.53, 0.57]
E 0.23 [0.22, 0.25] 0.52 [0.50, 0.53] 0.38 [0.36, 0.40]
R 0.34 [0.32, 0.36] 0.20 [0.19, 0.21] 0.39 [0.37, 0.41]
Table 2: Bias score (point, 95% CI) by approach and model. BBQ results obtained from Gallegos et al. (2025); CIs were only provided for GPT-3.5-Turbo. Mitigation techniques are Explanation (E), Reprompting (R) compared to Baseline (B).
The results reveal two key findings regarding the LACES and BBQ benchmarks. First, the baseline bias score (before any debiasing) is significantly higher for LACES, averaging 4.92 times that of BBQ (for example, 0.53 against 0.12 for GPT-3.5-Turbo). This finding reveals that the debiasing potential on LACES is much higher than on BBQ. Second, both tested debiasing approaches are less effective on LACES: explanation (E) reduces the baseline bias by 89% on BBQ but only 37% on LACES, whereas re-prompting (R) achieves an 80% reduction on BBQ compared to 45% reduction on LACES. LACES poses a greater challenge to current debiasing methods.
@@ 6 Discussion and Conclusions
Over time, concerted effort has been made towards the intentional growth of stereotype resources to be representative of global populations and perspectives. While the cost and scale of such collection remains a tradeoff, our approach to data collection builds a unique blend of techniques that eases this balance. Our methodology introduces an adaptive approach that is able to dynamically select data points that the annotator is able to validate and pivot through to extend the dataset. This is particularly important as it is able to situate the annotator in a particular topic and prevents a reliance on de-contextualized and static examples selected by the researchers. Section 5.2 shows the extent of unique concepts uncovered as compared to other similar datasets.
With this, we share a dataset of 4789 stereotypes, covering 120 identities and 842 attributes, as annotated by 83 individuals from 15 distinct countries, primarily from Latin-America, thus successfully covering perspectives not commonly covered in popular stereotype resources in NLP. The dataset’s diverse topics, detailed in Section 4, range from the widely accepted to the highly controversial, as identified through the participants reported recognition of the stereotypes.
This dataset not only includes a variety of new associations, but also shows the importance of multiple points of view. In particular, by studying inter-annotator agreement, we highlight how annotators portray a bias when validating stereotypes about themselves. They agree more with positive and neutral attributes, when it is about their group, than when it is about another group. Finally, we demonstrate how our resource can play a critical role in assessing self-debiasing techniques. These techniques proved successful with other data but work significantly worse in ours, suggesting that stereotypes covered in LACES are hard to debias.
@@ Appendix A Annotators and Validations
In Figure 7 we can observe the number of interactions across annotator nationality. Latin American annotations make up the vast majority of the data, with the exception of Spain.
@@
Figure 8: Number of validations per data point. Most data points were annotated only once.

### v2

@@ 5.1 Unique Attributes
To quantify the conceptual novelty of LACES, we measure its thematic overlap with stablished and recent benchmarks: CrowS-Pairs (Nangia et al., 2020), BBQ (Parrish et al., 2022), SeeGULL (Jha et al., 2023), SHADES (Mitchell et al., 2025), and HESEIA (Ivetta et al., 2025).
Our methodology employs a semantic vector space approach to identify thematic redundancies rather than relying on exact keyword matching. We transformed each data point into a high-dimensional representation using the multilingual text-embedding-3-large model (OpenAI, 2026). For every entry across all datasets, we calculated the maximum cosine similarity against all entries in the comparison pool. An attribute is classified as "unique" if its embedding does not exceed a predefined similarity threshold with any existing record in the other datasets, indicating that the entry represents a novel thematic contribution.
As shown in Table 2, the results demonstrate the percentage of unique concepts within each dataset when analyzed for thematic overlap, revealing that the LACES dataset contains the highest percentage of unique concepts at 29.74%. This figure represents an improvement over other regional and global benchmarks, surpassing HESEIA (27.78%) and SeeGULL (19.63%), and significantly over the uniqueness found in other resources such as BBQ (15.35%), CrowS-Pairs (14.33%), and SHADES (10.85%). The overall conceptual overlap between datasets can be visualized in Appendix C.
Dataset Size Unique Attributes
LACES 4,879 1,650 29.74%
HESEIA 45,416 13,134 27.78%
SeeGULL 25,861 6,716 19.63%
BBQ 58,492 11,452 15.35%
CrowS-Pairs 1,508 321 14.33%
SHADES 728 79 10.85%
Table 2: Comparison of Dataset Size and Attribute Diversity. This table presents size, total unique attributes, and the percentage of unique concepts (highest bolded) across evaluated datasets. The uniqueness percentage is calculated using thematic similarity embeddings.
Interestingly, while both participatory benchmarks (HESEIA and LACES) contain the highest proportions of unique concepts, LACES proves more efficient. HESEIA took approximately 8,535 hours to collect (1.5 hours in average per 5,690 participants) while LACES took approximately 300 hours (3.6 hours in average per 83 participants). Despite LACES containing ten times fewer data points than HESEIA and only a small fraction of participants, it yields a higher percentage of unique entries. We hypothesize that this is a result of our adaptive methodology, which actively incentivizes annotators to explore the "long tail" of sociocultural knowledge. In contrast, the static examples provided in HESEIA and other resources, in the annotation guidelines, may have inadvertently constrained annotator exploration to more salient attributes.
@@ 5.2 Self-Debiasing
To evaluate the generalizability of zero-shot self-debiasing, we extend techniques proven effective in BBQ to the datasets analyzed in the previous section. HESEIA was excluded as it does not distinguish between stereo- and anti-stereotypes, which is required for this experiment. We adopt the BBQ evaluation protocol, using multiple-choice items that ask which country is associated with a specific stereotypical attribute, including an "Unknown" option to mitigate forced-choice bias. Following Gallegos et al. (2025), we evaluate three debias strategies (see Appendix B for details):
Baseline (B): A direct prompt for a single-letter answer without reasoning or intervention.
Explanation (E): An approach where the model identifies invalid assumptions before answering.
Reprompting (R): A two-step intervention instructing to remove bias from its initial response.
Parrish et al. (2022) defines $\textsc{Bias}=(1-\textsc{ACC})\left[2\frac{n_{\textsc{biased}}}{m}-1\right]$, ACC is accuracy, $m$ is the number of non-Unknown predictions, and $n_{\textsc{biased}}$ counts predictions matching the bias target.
Dataset S GPT Llama-3.1
3.5-Turbo 4o-mini
LACES B 0.53 0.68 0.55
E 0.23 0.52 0.38
R 0.34 0.20 0.39
SeeGULL B 0.27 0.31 0.24
E 0.09 0.15 0.13
R 0.12 0.02 0.11
BBQ B 0.12 0.14 0.10
E 0.03 0.01 0.00
R 0.04 0.01 0.02
CrowS-Pairs B 0.18 0.15 0.18
E 0.04 0.01 0.03
R 0.09 0.00 0.07
SHADES B 0.34 0.23 0.15
E 0.04 0.07 0.10
R 0.10 0.00 0.15
Table 3: Bias score by benchmark (the larger the number, the more biased). Debiasing strategies (S) include Explanation (E), Reprompting (R), compared to Baseline (B) following BBQ methdology Gallegos et al. (2025). Highest bias scores per model/strategy are bold.
As shown in Table 3, LACES exhibits the highest average baseline bias across all benchmarks, significantly exceeding the runner-up: SeeGULL. Notably, SeeGULL was generated by a diverse pool of LLMs and validated through approximately 1,155 hours of human validation from all over the world. It represents a high-quality, best-case LLM-driven benchmark.
Current debiasing techniques suffer from a marked performance decay when applied to our benchmark. The Explanation (E) strategy achieves an average bias reduction of 66% across datasets, ranging from 89% in BBQ to 55% in SeeGULL, but only reaches 37% in LACES. Reprompting (R) reduces bias by an average of 64% across benchmarks (including 80% in BBQ and 70% in CrowS-Pairs), but its effectiveness falls to 45% for LACES.
Collectively, these results suggest that standard mitigation approaches are brittle when applied to regional nuances and sociocultural knowledge, highlighting a critical gap in fairness research.
@@ 6 Discussion and Conclusions
Over time, concerted effort has been made towards the growth of stereotype resources to be representative of global populations and perspectives. Our proposed methodology integrates adaptive sampling techniques to mitigate the inherent trade-offs between collection scale, cost and coverage. By dynamically prioritizing data points for validation and facilitating community-driven pivots, the framework situates annotators within relevant sociocultural contexts, thereby reducing the reliance on decontextualized seed examples. Section 5.1 explores how this approach surfaces a high proportion of unique concepts absent from similar benchmarks.
With this, we introduce LACES: a dataset of 4789 stereotypes, covering 120 identities and 842 attributes, as annotated by 83 individuals from 15 distinct countries, primarily from Latin-America, thus successfully covering perspectives not commonly covered in popular stereotype resources in NLP. The dataset’s diverse topics, detailed in Section 4, range from the widely accepted to the highly controversial, as identified through the participants reported recognition of the stereotypes.
The efficiency of our adaptive methodology is further evidenced by a comparative analysis with the HESEIA dataset Ivetta et al. (2025). While HESEIA represents a significant effort in documenting Latin American social biases through a large-scale pedagogical initiative, LACES yields a higher proportion of unique associations with approximately 10% of the data volume. This disparity suggests that our adaptive sampling algorithm and task design effectively optimize the data collection process, reducing resource expenditure without sacrificing diversity. While HESEIA serves broader societal objectives such as teacher training, LACES offers a high-precision alternative for achieving data saturation with minimal overhead.
The architecture of the framework provides a scalable foundation by allowing future collections to customize the sampling algorithm and weight distributions to align with specific research priorities. This flexibility enables the prioritization of different data collection strategies, such as maximizing geographic coverage, centering specific minority perspectives, or increasing validation counts for data points with high inter-annotator disagreement to improve reliability, among others.
This transition points toward a broader paradigm shift in how fairness resources are curated, moving from static repositories to interactive workflows. This may have implications for sociolinguistic methodologies that consider human language as proposed in Plank (2022b). This participatory approach encourages that captured associations are grounded in authentic community insights rather than being constrained by researcher bias. By shifting epistemic authority to grassroots communities, the LACES framework allows datasets to function as dynamic resources that adapt to evolving sociocultural contexts.
@@ Appendix A Annotators and Validations
In this section we describe basic validation and annotation data. In Figure 5 we can observe the number of interactions across annotator nationality. Latin American annotations make up the vast majority of the data, with the exception of Spain.
@@
Figure 6: Number of validations per data point. Most data points were validated once.
@@ Appendix B Self-Debiasing
This section includes details about the methodology to unify dataset format so that the self-mitigation strategies could be applied. The methodology requires the datasets to distinguish stereotypes from anti-stereotypes. Also, the promptins strategies are included for reproducibility
@@ B.3 Evaluation
Appendix C Attributes Overlap among Benchmarks
Figure 8 provides a visualization of the conceptual overlap among the datasets. The width of the chords connecting the nodes quantifies the number of shared concepts between any two datasets, while the color of the chord indicates the dataset where the concept originated. The most notable insight is that the LACES dataset is the primary source of unique concepts, indicated by its large unconnected arc segment and the least overlap with other datasets. Furthermore, SHADES exhibits the largest mutual conceptual overlap across the board, as demonstrated by the thickest connecting chords, suggesting it addresses underlying concepts already explored in other resources. Interestingly, not all conceptual overlaps are symmetric; for example, a large number of concepts from BBQ can be found in HESEIA but less than half that amount can be found the other way round.